OpenStudy (elleblythe):

If a rowing team rows with the current, they can cover 16 miles in 2 hours. Against the current, they can cover 10 miles in 4 hours. Find their rate of rowing in still water and the rate of the current. Show solution please if possible

OpenStudy (anonymous):

put the rate of the current as say (r\) and the rate in still water as \(x\) then since distance is rate times time you know \[16=2(x+r)\] and also \[10=4(x-r)\] solve this for \(x\) and \(r\)